# VQE for Ising Model & A Comparative Analysis of Classical and Quantum Optimization Methods

### v1

@@ V.1 Complexity estimation
Derivative-based method efficiency $\triangledown f({\bm{\theta}})$ PSR FD SPSA Comp. cost 2p 2p 2 $g({\bm{\theta}})$ HES-PSR QN-BDA QN-SPSA Comp. cost 3p.p L 4
Table 1: Showing comparison among combinations of gradient and adaptive learning rate methods. Where p is the number of parameters, L is the number of PQC layers of the same rotation generator.
@@
In summary of the derivatives-based optimization framework, we tell apart two subroutines: gradient calculation and $g({\bm{\theta}})$ matrix evaluation. For gradient calculation, we got three approaches: PSR offering exact gradients of quantum ansatz structure, FD-a classical numerical method that approximates gradients, and SPSA-a stochastic approximation technique requiring only two function evaluations per iteration. For the $g({\bm{\theta}})$ matrix calculation, we generally have four methods: Const—an identity matrix; HES-PSR, which provides an exact second-order Hessian matrix by the PSR method of the quantum structure objective; QN-BDA approximates the Fubini-Study metric using the block diagonal structure; and QN-SPSA, which stochastically approximates the FIM with only four evaluations per iteration. A method is a combination of the gradient and $g({\bm{\theta}})$, such as QN-BDA+PSR, which implies the application of the QN-BDA technique and PSR for the estimation of the gradient.
Figure 3: The illustratively comparative graphics among combinations of gradient and adaptive learning rate methods.
@@
We provide a rough assessment of the optimization methods, grounded in theoretical judgment, where you are able to read off their information from Tab. 1 and Fig. 3 for more illustrative purpose, demonstrating the trade-offs in performance and resource requirements among methods, e.g., computational cost, accuracy, and convergence speed. Tab. 1 is derived from the underlying theories of methods discussed above. Besides that, Fig. 3 provides an intuitive estimation based on theory-driven and bias-informed conjectures. In this estimation, the methods HES-PSR, QN-BDA, QN-SPSA, and Const are ranked in descending order of convergence, accuracy, and computational cost. Similarly, for gradient methods, SPSA, FD, and PSR follow an ascending order of accuracy, convergence, and computational cost. Consequently, one can infer, for instance, that HES-PSR+PSR should achieve the best performance at the expense of the intractably highest computational cost, while QN-BDA+PSR consumes much quantum overhead but is expected to outperform QN-SPSA+SPSA overall.
@@
From this framework, a key motivation of the QN-SPSA+PSR algorithm is to merge the computational resource efficiency of QN-SPSA with the precision of the PSR. The computational cost of QN-SPSA, at four quantum evaluations per iteration, demonstrates remarkable scalability compared to traditional Hessian-based approaches like HES-PSR or QN-BDA, which scale quadratically and linearly with the number of parameters and layers, while the PSR would keep the method at good stability and convergence. This refinement enables faster convergence and greater stability while maintaining a low computational overhead, highlighting its potential for practical application on NISQ devices. Notably, these are only theory-based rough estimations. In practice, the complexity of objective functions and the influence of noise would lead to variations in the performance of these methods, and approximate methods, such as methods using SPSA, would have advantages as noises come about. The experimentally numerical results, however, demonstrate an even better performance of QN-SPSA+PSR than predicted by the estimation, as we will present in the next section.
(a) Linear entanglement mapping.
(b) Full entanglement mapping.
Figure 4: Comparing multiple optimizations of TIM’s 12 spins with the RealAmplitudes Ansatz.
V.2 Simulation results
@@
Figure 5: Comparing multiple optimizations of TIM’s 12 spins with a) the RealAmplitudes Ansatz and b) the EfficientSU2 Ansatz.
@@
Furthermore, Figs. 4 and 5, which compared two types of ansatz and two types of entanglement schemes, turned out to show that linear entanglement is not much different than the full entanglement, and the RealAmp performed better in convergence than Efficient SU2. That proved the initial arguments aforementioned in section III.2.2, the RealAmplitude with linear entanglement, therefore, is good enough to survey the Ising model. Finally, the TIM energy estimates are obtained by varying external fields and numbers of spins, as illustrated in Figs. 6 and 7. In both cases, the QN-SPSA+PSR consistently delivered reliable and accurate results for the average ground state energy, closely aligning with the QN-BDA+PSR.
VI Conclusion

### v2

@@ IV.1 Complexity estimation
Derivative-based method efficiency $\triangledown f({\bm{\theta}})$ PSR FD SPSA Comp. cost 2p 2p 2 $g({\bm{\theta}})$ QN-BDA QN-SPSA Const Comp. cost L 4 0
Table 1: Comparison of derivative-based optimization methods in terms of computational cost. The table presents different optimization algorithms with their associated computational costs, respectively. The methods included are Parameter Shift Rule (PSR), Finite Differences (FD), and Simultaneous Perturbation Stochastic Approximation (SPSA), referring to gradient estimation techniques; Quantum Natural Gradient with Block-Diagonal Approximation (QN-BDA), Quantum Natural Gradient with SPSA (QN-SPSA), and a constant learning rate (Const) for adaptive learning rate methods; where p is the number of parameters and L is the number of PQC layers of the same rotation generator.
@@
In summary of the derivatives-based optimization framework, we tell apart two subroutines: gradient calculation and $g({\bm{\theta}})$ matrix evaluation. For gradient calculation, we got three approaches: PSR offering exact gradients of quantum ansatz structure, FD-a classical numerical method that approximates gradients, and SPSA-a stochastic approximation technique requiring only two function evaluations per iteration. For the $g({\bm{\theta}})$ matrix calculation, we generally have four methods: Const—an identity matrix; QN-BDA approximates the Fubini-Study metric using the block diagonal structure; and QN-SPSA, which stochastically approximates the Fisher Information Matrix (FIM) with only four evaluations per iteration. A method is a combination of the gradient and $g({\bm{\theta}})$, such as QN-BDA+PSR, which implies the application of the QN-BDA technique and PSR for the estimation of the gradient.
Figure 3: The illustratively comparative graphics among combinations of gradient and adaptive learning rate methods in terms of accuracy and convergence speed. The figure illustrates the performance of different optimization strategies, categorized by gradient estimation methods (PSR, FD, SPSA) and adaptive learning rate techniques (QN-BDA, QN-SPSA, Const). The results presented here are derived from the findings of previous works [45, 46, 50, 51, 52], alongside theoretical predictions and experimentally numerical scans conducted in this study.
@@
We provide a rough assessment of the optimization methods, grounded in theoretical judgment, where you are able to read off their information from Tab. 1 and Fig. 3 for more illustrative purpose, demonstrating the trade-offs in performance and resource requirements among methods, e.g., computational cost, accuracy, and convergence speed. Tab. 1 is derived from the underlying theories of methods discussed above. Besides that, Fig. 3 provides an intuitive estimation based on theory-driven and bias-informed conjectures. In this estimation, the methods QN-BDA, QN-SPSA, and Const are ranked in descending order of convergence, accuracy, and computational cost. Similarly, for gradient methods, SPSA, FD, and PSR follow an ascending order of accuracy, convergence, and computational cost. Consequently, one can infer, for instance, that while the fully computed Hessian-like matrix+PSR should theoretically achieve the best performance at the expense of the intractably highest computational cost, the feasible approximate QN-BDA+PSR consumes much quantum overhead but is expected to outperform QN-SPSA+SPSA overall.
@@
From this framework, a key motivation of the QN-SPSA+PSR algorithm is to merge the computational resource efficiency of QN-SPSA with the precision of the PSR. The computational cost of QN-SPSA, at four quantum evaluations per iteration, demonstrates remarkable scalability compared to traditional Hessian-based approaches like QN-BDA, which scale quadratically or linearly with the number of parameters and layers. While the PSR would keep the method at good stability and convergence. This refinement enables faster convergence and greater stability while maintaining a low computational overhead, highlighting its potential for practical application on NISQ devices. Notably, these are only theory-based rough estimations. In practice, the complexity of objective functions and the influence of noise would lead to variations in the performance of these methods, and approximate methods, such as methods using SPSA, would have advantages as noises come about. The experimentally numerical results, however, demonstrate an even better performance of QN-SPSA+PSR than predicted by the estimation, as we will present in the next section.
Figure 4: Comparison of optimization methods for the TIM model with 12 spins using the RealAmplitudes ansatz. The plots show the relative error as a function of iterations for different optimization strategies applied to two entanglement mappings: (a) linear entanglement mapping and (b) full entanglement mapping. The plot illustrates the convergence behavior of each method, also showing the comparable results of the linear entanglement scheme to the full scheme, yet at a low quantum cost.
IV.2 Simulation results
@@
Figure 5: Comparison of optimization methods for the TIM model with 12 spins using different ansatz configurations. The plots illustrate the relative error as a function of iterations for multiple optimization strategies applied to (a) the RealAmplitudes ansatz and (b) the EfficientSU2 ansatz. The visualized outcomes highlight the better and compatible performance of the RealAmplitudes ansatz for the TIM problem, as previously discussed in Section II.3.2.
V Conclusion
Furthermore, Figs. 4 and 5, which compared two types of ansatz and two types of entanglement schemes, turned out to show that linear entanglement is not much different than the full entanglement, and the RealAmp performed better in convergence than Efficient SU2. That proved the initial arguments aforementioned in Section II.3.2, the RealAmplitude with linear entanglement, therefore, is good enough to survey the Ising model. Finally, the TIM energy estimates are obtained by varying external fields and numbers of spins, as illustrated in Figs. 6 and 7. In both cases, the QN-SPSA+PSR consistently delivered reliable and accurate results for the average ground state energy, closely aligning with the QN-BDA+PSR.
@@ V Conclusion
Acknowledgements.
The authors acknowledge helpful discussions and contributions from Triet Minh Ha. The numerical calculations in this work are performed using the high-performance computing systems at Phenikaa University. This work is financially supported by the Ministry of Education and Training of Vietnam under the grant number QG.20.17.
Author contributions: V.D. Nguyen and H.Q. Nguyen guided and supervised the project. The ideas and results were developed and carried out by D.T. Le and V.L. Nguyen, with support and discussions from V.D. Nguyen, H.Q. Nguyen, and H.C. Nguyen. D.T. Le wrote the manuscript with assistance from V.L. Nguyen and feedback from H.Q. Nguyen, V.D. Nguyen, and H.C. Nguyen. V.L. Nguyen packaged the code and executed the final summarization results on the HPC system.
Data availability: The datasets generated and analysed during the current study are available in the Variational-Quantum-EigeinSolver repository, https://github.com/nguyenvulinh666/Variational-Quantum-EigeinSolver.
Declarations
Competing interests: The authors declare no competing interests.